### v1

@@ 2.2 Linear equation, bounded disc
Similarly to the case of a freely expanding disc, a method of superposition of particular solutions to Eq. (3) can be applied. The difference is that, for a bounded disc with $r_{\mathrm{out}}=$const, the corresponding eigenvalue problem has a purely discrete spectrum. Therefore, the Green’s function is not an integral, but a linear combination of particular solutions satisfying the specific boundary conditions. A particular Green’s function for a finite disc was constructed by Wood et al. [53]. The full Green’s function, which can be used for an arbitrary initial density distribution for the two types of inner boundary conditions, has been constructed by Lipunova [48] for $r_{\mathrm{in}}=0$; see also Mushtukov et al. [54] for $r_{\mathrm{in}}>0$.
Table 1: Parameters of the time-dependent solution with index $a=0$ in viscosity $\nu\propto\Sigma^{a}\,r^{b}$. The ring starts at $r_{\rm s}$ and evolves into the disc with zero inner torque. If the disc expands freely then the central accretion rate decays as $\propto t^{-1-l}$, see Eq. (10). If the disc has a fixed outer boundary, $\dot{M}$ decays exponentially, see Eq. (16). For such disc the peak time is similar to the case of a freely expanding case  (13), for which it is the exact value. Parameter $k_{1}$ is an eigenvalue for the solution in the interval ($0,r_{\rm out}$). Solutions with $a=0$ and specific $b$ approximate the $\alpha$-disc evolution on time scales of the order of a couple of viscous timescales. Different opacity laws (the Rosseland mean) are indicated for $\alpha$-discs.
@@
Disc with $z/r=$const, e.g., ADAF $1/2$ $1/3$ 1.24 $1/9$ $0.38$
@@
where $\nu_{\mathrm{out}}=\nu_{0}\,r_{\mathrm{out}}^{b}$ and $12\,l\,(2-b)=$ 1.
In Table 1 some parameters of solutions with $m=0$ are presented, when the outer boundary condition $\dot{M}(r_{\mathrm{out}})={\partial F}/{\partial h}=0$ and the inner boundary condition $F(r_{\mathrm{in}})=0$. The peak time for a bounded disc is of the order of $t_{\rm max}$ for a freely expanding disc, but it also depends on the outer radius $r_{\rm out}$ [48]. The disc becomes quasi-stationary (namely, the accretion rate virtually does not change with radius) in the region limited by $r<r_{\mathrm{out}}\times(t/t_{\mathrm{exp}})^{2\,l}$. The establishment of quasi stationarity in the central regions of discs on viscous time scales is a common property for discs with any type of viscosity.
Nixon and Pringle [41] found the Green’s function for $\nu\propto r$ and a general inner boundary condition (14). They assume the outer boundary condition $F(r_{\mathrm{out}})=0$ or $\Sigma(r_{\mathrm{out}})=0$, so that the mass and angular momentum reaching the outer boundary is absorbed there.
Some bright outbursts in X-ray novae show lightcurves with fast rise and exponential decay, which are called FRED profiles [55]. King and Ritter [56] studied the evolution of a disc with the finite outer radius and constant $\nu_{\mathrm{t}}$, and found that the accretion rate declined exponentially with time. This does not necessarily mean that the viscosity is constant in the real discs. The FRED light curves are equally well approximated within the model of evolving $\alpha$-discs with a time-dependent viscosity $\nu$. This is due to the fact that (i) on time scales of the order of one to two $t_{\mathrm{vis}}$, an $\alpha$-disc and a disc with time-independent viscosity show similar evolution (compare the solid curves in Fig. 1); (ii) the typical spectral evolution of an X-ray Nova in the soft state implies a softening of the spectrum, characterized then by an exponential tail. Consequently, a power-law variation of the disc maximum temperature is masked by the exponentially decreasing flux in the observational band [57]. In any case, before applying an analytic model to an outburst of an X-ray nova, it is necessary to ascertain whether the radius of the high-viscosity zone in the disc is constant or the cooling front is propagating inwards, see §4.
@@
Stationary $\alpha$-discs with the dominant gas pressure and Kramers opacity have the aspect ratio $z/r\propto r^{1/8}\dot{M}^{3/20}$ [1]. Thus $b=3/4$ if we ignore the dependence of the disc thickness on the accretion rate. Further options for $\alpha$-discs are listed in Table 1. Potentially, knowing the observed time $t_{\exp}$, one can estimate $\alpha$ in an X-ray nova using (16) and (17[48, 58]. However, the main uncertainty arises from the unknown outer radius of the high-viscosity zone. The most promising approach to determine $\alpha$ observationally lies in examining the FRED outbursts of the short-period BH XRBs, where there is a high chance that $r_{\rm out}$ coincides with the outer radius of a tidally truncated disc.
@@ 2.3 Solutions to nonlinear evolution equation for freely expanding discs
Earlier we considered a scenario in which the kinematic viscosity coefficient depends only on the radial coordinate inside the disc. For many viscosity mechanisms, we can express $\nu$ as a function of a power law of $\Sigma$ and $r$. Consequently, the viscous time in the disc varies with the radius and time. This is relevant, for example, for discs with $\alpha$-viscosity. For solving the nonlinear differential equation of disc evolution (3), similarity methods have proven to be quite effective.
Self-similar solutions to nonlinear differential equations can be divided into two kinds [59, 60]. The self-similar solution is of the first kind if the self-similar function, as well as its dimensionless argument, can be derived from dimensional analysis or conservation laws. This case is also called a complete self-similarity. This is, for example, J. I. Taylor’s blast wave problem with energy conservation.
The second kind, or incomplete self-similarity, arises when it is not possible to use dimensional analysis to determine the self-similar function, or to find the powers to which the dimensional parameters should be raised to produce a self-similar dimensionless variable. In this case, the self-similar function is found as a particular solution to the problem itself [a ‘nonlinear eigenvalue problem’, 61]. The self-similar solution describes the ‘intermediate asymptotic’ behavior of the system in the region where it no longer depends on the initial and boundary conditions.
Self-similar solutions of the first kind have been found by Pringle [62, 39] for the nonlinear viscous diffusion equation of accretion discs in the evolutionary stage when the accretion rate decays. The conservation of total angular momentum in an accretion disc is used there.
Solutions of the second kind have been obtained by Lyubarskij and Shakura [37] for the earlier evolutionary stage when the central accretion rate increases.
Lyubarskij and Shakura [37] proposed a division of the whole evolution of a ring into three stages, each of which allows self-similar solutions: (1) the stage of initial spreading of the disc from a ring, (2) the gradual development of a quasi-stationary distribution of parameters in the disc, and (3) the late ‘spreading’ of the disc away from the center, accompanied by a decrease in the central accretion rate.
@@
The disc gradually evolves into the third and final stage (the decay stage), when the details of the initial distribution are ‘forgotten’ anyway, and only some integral quantities conserved by the evolution are important in finding the self-similar solution. [^3] The nonlinear problem has the following distinctive features. Firstly, the self-similar solutions of the second kind exist only for $m\neq 0$. Secondly, self-similar solutions of the first kind in the third stage, while they exist for $m=0$, have an exponential radial profile for $r\rightarrow\infty$, characteristic for linear problems [see, for example, 29]. For $m\neq 0$, the position of the boundary of the disc is fully determined, see also Rafikov [40]. This property is similar to the one that arises in problems of thermal conductivity, when, due to the nonlinearity, the heatwave boundary sharply separates the heated zone from the rest of the region [61]. The solution is sought in the form $F={h^{({n+2})/{m}}}\,{(\pm D\,t)^{-1/m}}\,y(\xi)$ where $y$ is a dimensionless function of a dimensionless argument $\xi={h}/({A\,(\pm t)^{\beta}})$. The index $\beta$ can be determined in the course of the solution for the first and second stage, and from the conservation of the angular momentum, for the decay stage (then $\beta$ can be expressed algebraically, see also Pringle 62). The accretion rate on the gravitating center decays as follows [see., e.g., chapter 1 in 31] [^4] For $a=0$ and $m=0$, (18) transforms into the Lynden-Bell and Pringle solution (10) at large $t$. :
@@
$\alpha-$disc, $\kappa=$ const, $P_{\rm tot}=P_{\rm rad}$2 3/2 (-8/9) $2$ $7$
@@
$\alpha-$disc, $\kappa\propto\rho T^{-5/2}$, $P_{\rm tot}=P_{\rm gas}$ $1/2$ $1$ $11/9$ $1/3$ $1$ $3.319$
convective turbulence, $\kappa\propto T^{2}$, Lin and Bodenheimer [63] 2 0 $15/14$ $2/3$ $8/3$ 4.820
molecular disc with gravitational instability, Lin and Pringle [30] 2 $9/2$ $6/5$ $2/3$ $-1/3$ 1.788
In the gas-pressure dominated $\alpha$-disc with the opacity according to the Kramers law and Thomson scattering, the accretion rate decreases as $\propto t^{-19/16}$ and $\propto t^{-5/4}$, respectively [62, 36, 37, 64, 39]. Actually, opacity $\kappa\propto\rho T^{-5/2}$ better describes $\alpha$-disc around stellar-mass objects for conditions near the outer radius of the hot ionized zone [65]. Other solutions in Table 2 are important in the context of the protoplanetary discs and discs in galactic nuclei. Lin and Pringle [30] considered a molecular disc with a gravitational instability producing an effective viscosity. A different approach to describe analytically viscosity in self-gravitating discs, depending on their thermal structure and external irradiation, can be found in Rafikov [66, 67]. Lin and Bodenheimer [63] studied the evolution of a protoplanetary disc under the influence of convective turbulent viscosity.
Pringle [39] considered an “external” disc with a zero central accretion rate and a finite central torque, w [a massive binary surrounded by a gas disc, see also 68]. For all mass initially at the origin and being gradually expelled to infinity, the inner torque decays as
@@
The latter dependence converges to the solution for the linear equation with $a=0$, see §2.1. In Rafikov [38], self-similar solutions were found with a non-zero constant mass accretion rate at the inner boundary $\dot{M}(R_{\rm in})=\chi\dot{M}_{\infty}$ and a constant external mass supply $\dot{M}_{\infty}$. Larger $\chi$ values imply less mass accumulation in the disc and a smaller inner torque. General late-time power-law asymptotics for freely expanding discs with simultaneously non-zero central accretion rate and non-zero central viscous torque at the inner boundary were presented by Rafikov [40]. Formulating such solutions required setting the inner coordinate of the disc to zero.
A time-dependent solution of Ogilvie [69] for accretion flows that retain the heat generated by the viscous dissipation — ’advection-dominated’ flows – is another example of using the similarity methods. In the case of quasi-spherical accretion, the governing equations (which are spherically averaged) are nonlinear, so that the self-similar solution of the first kind is an attractor for the initial-value problem with the conserved angular momentum. The accretion rate varies $\propto t^{-4/3}$, which is consistent with the solution of the thin-disc evolution equation with $z/r=$const, see Table 1.
@@ 2.4 Non-linear evolution equation for $r_{\mathrm{out}}=\,$const
Figure 1: Comparison of analytic solutions (12), (16), (18), and (20) for the evolution of the accretion rate onto a 10 $M_{\odot}$ black hole with different viscosity and disc outer boundary (freely expanding or fixed). The viscosity in each solution is chosen to approximate the law in the zone of dominant gas pressure and opacity according to the Kramers formula. At late time, the time-dependencies tend to $(t/\tau_{\mathrm{pl}})^{-7/5}$, $\exp(-t/t_{\mathrm{exp}})$, $(1+t/\tilde{t}_{0})^{-5/4}$, and $(1+t/t_{0})^{-10/3}$, respectively. The input parameters are $M_{\rm disc}=6\times 10^{24}~{}$g, $R_{\mathrm{out}}=2\times 10^{11}$ cm and $t_{\rm vis}=60^{\mathrm{d}}$. The corresponding $e$-decay time is $t_{\mathrm{exp}}\approx 27^{\mathrm{d}}$; the bounded $\alpha$-disc’s solution is close to the linear-problem solution for about $3\,t_{\mathrm{exp}}$.
A solution to the basic equation of non-stationary accretion (3) for a disc with a constant outer radius can be found during the decay stage using the method of separation of variables: $F(h,t)=F(t)\times f(h/h_{0})$. The accretion rate changes proportionally to $F(t)$ [70, 57, 71]:
@@
where $\dot{M}_{0}$ is the central accretion rate at $t=0$, which can be chosen as any time at the decay stage. The space part of the solution $f(h/h_{0})$ for a disc with the zero (or very small) viscous stress at the inner boundary was found semi-analytically as a polynomial by Lipunova and Shakura [57]. The outer boundary condition was set to $\dot{M}(r_{\mathrm{out}})=0$, which is relevant during outbursts of X-ray novae, when the accretion rate inside the disc is much higher than the mass transfer rate from the companion star. The full solution allows determination of $t_{0}$, which is of the order of the viscous time at the disc outer radius:
@@
and the time scale on which the disc mass decays initially [71]. Here $h_{\rm out}=\sqrt{GMr_{\rm out}}$, $a_{0}\equiv f^{\prime}(0)$, $a=m/(1-m)$, and $\lambda\sim 3$ (see Table 2). The solution for the Kramers opacity was used to model the optical and X-ray lightcurves of the X-ray novae A 0620-00 and GU Mus 1124-68 during the decline after the peak of their outbursts and to obtain constraints on the turbulence parameter $\alpha$ [72, 73].

### v2

@@ 2.2 Linear equation, bounded disc
Similarly to the case of a freely expanding disc, the method of superposition of particular solutions to Eq. (3) can be applied. The difference is that, for a bounded disc with $r_{\mathrm{out}}=$const, the corresponding eigenvalue problem has a purely discrete spectrum. Therefore, the Green’s function is not an integral, but a linear combination of particular solutions satisfying the specific boundary conditions. A particular Green’s function for a finite disc was constructed by wood_etal2001. The full Green’s function, which can be used for an arbitrary initial density distribution for the two types of inner boundary conditions, has been constructed by lipunova2015 for $r_{\mathrm{in}}=0$ (Fig. 1, the right panel); see also Mushtukov+2019 for $r_{\mathrm{in}}>0$.
Table 1: Parameters of the time-dependent solution with index $a=0$ in viscosity $\nu\propto\Sigma^{a}\,r^{b}$. The ring starts at $r_{\rm s}$ and evolves into the disc with zero inner torque. If the disc expands freely then the central accretion rate decays as $\propto t^{-1-l}$, see Eq. (10). If the disc has a fixed outer boundary, $\dot{M}$ decays exponentially, see Eq. (16). For such a disc the peak time is similar to the case of a freely expanding case  (13), for which it is the exact value. Parameter $k_{1}$ is an eigenvalue for the solution in the interval ($0,r_{\rm out}$). Solutions with $a=0$ and specific $b$ approximate the $\alpha$-disc evolution on time scales of the order of a couple of viscous timescales. Different opacity laws (the Rosseland mean) are indicated for $\alpha$-discs.
@@
Disc with z/r=z/r=const, e.g., ADAF $1/2$ $1/3$ 1.24 $1/9$ $0.38$
@@
where $\nu_{\mathrm{out}}=\nu_{0}\,r_{\mathrm{out}}^{b}$ and 2l(2b)=12\,l\,(2-b)=1.
In Table 1 some parameters of solutions with $m=0$ are presented, when the outer boundary condition $\dot{M}(r_{\mathrm{out}})={\partial F}/{\partial h}=0$ and the inner boundary condition $F(r_{\mathrm{in}})=0$. The peak time for a bounded disc is of the order of $t_{\rm max}$ for a freely expanding disc, but it also depends on the outer radius $r_{\rm out}$ [lipunova2015]. The disc becomes quasi-stationary (namely, the accretion rate virtually does not change with radius) in the region limited by $r<r_{\mathrm{out}}\times(t/t_{\mathrm{exp}})^{2\,l}$. The establishment of quasi stationarity in the central regions of discs on viscous time scales is a common property for discs with any type of viscosity.
Nixon-Pringle2021 found the Green’s function for $\nu\propto r$ and a general inner boundary condition (14). They assume the outer boundary condition $F(r_{\mathrm{out}})=0$ or $\Sigma(r_{\mathrm{out}})=0$, so that the mass and angular momentum reaching the outer boundary is absorbed there.
Some bright outbursts in X-ray novae show lightcurves with fast rise and exponential decay [FRED lightcurves, chen_et1997]. kin-rit1998 studied the evolution of a disc with the finite outer radius and constant $\nu$, and found that the accretion rate declined exponentially with time. This does not necessarily mean that the viscosity is constant in real discs. The FRED light curves are equally well approximated within the model of evolving $\alpha$-discs with a time-dependent viscosity $\nu$. This is due to the fact that (i) on time scales of the order of one to two $t_{\mathrm{vis}}$, an $\alpha$-disc and a disc with time-independent viscosity show similar evolution (compare the solid curves in Fig. 3); (ii) the typical spectral evolution of an X-ray Nova in the soft state implies a softening of the spectrum, characterized then by an exponential tail. Consequently, a power-law variation of the disc maximum temperature is masked by the exponentially decreasing flux in the observational band [lipunova_shakura2000]. In any case, before applying an analytic model to an outburst of an X-ray nova, it is necessary to ascertain whether the radius of the high-viscosity zone in the disc is constant or the cooling front is propagating inwards, see §4.
@@
Stationary $\alpha$-discs which are gas-pressure dominated and have Kramers opacity have an aspect ratio $z/r\propto r^{1/8}\dot{M}^{3/20}$ [sha73]. Thus $b=3/4$ if we ignore the dependence of the disc thickness on the accretion rate. Further options for $\alpha$-discs are listed in Table 1. Potentially, knowing the observed time $t_{\exp}$, one can estimate $\alpha$ in an X-ray nova using (16) and (17[lipunova2015, lipunova-malanchev2017]. However, the main uncertainty arises from the unknown outer radius of the high-viscosity zone. The most promising approach to determine $\alpha$ observationally lies in examining the FRED outbursts of the short-period BH XRBs, where there is a high chance that $r_{\rm out}$ coincides with the outer radius of a tidally truncated disc.
@@ 2.3 Solutions to nonlinear evolution equation for freely expanding discs
Earlier we considered a scenario in which the kinematic viscosity coefficient depends only on the radial coordinate inside the disc. For many viscosity mechanisms, we can express $\nu$ as a function of a power law of $\Sigma$ and $r$. Consequently, the viscous time in the disc varies with radius and time. This is relevant, for example, for discs with $\alpha$-viscosity. For solving the nonlinear differential equation of disc evolution (3), similarity methods have proven to be quite effective.
Self-similar solutions to nonlinear differential equations can be divided into two kinds [barenblatt1982e, barenblatt2003scaling]. The self-similar solution is of the first kind if the self-similar function, as well as its dimensionless argument, can be derived from dimensional analysis or conservation laws. This case is also called a complete self-similarity. This is, for example, J. I. Taylor’s blast wave problem with energy conservation.
The second kind, or incomplete self-similarity, arises when it is not possible to use dimensional analysis to determine the self-similar function, or to find the powers to which the dimensional parameters should be raised to produce a self-similar dimensionless variable. In this case, the self-similar function is found as a particular solution to the problem itself [a ‘nonlinear eigenvalue problem’, zeldovich-raizer1967]. The self-similar solution describes the ‘intermediate asymptotic’ behavior of the system in the region where it no longer depends on the initial and boundary conditions.
Self-similar solutions of the first kind have been found by pringle1974PhD, pringle1991 for the nonlinear viscous diffusion equation of accretion discs in the evolutionary stage when the accretion rate decays. The conservation of total angular momentum in an accretion disc is used there.
Solutions of the second kind have been obtained by lyub-shak1987 for the earlier evolutionary stage when the central accretion rate increases.
lyub-shak1987 proposed a division of the whole evolution of a ring into three stages, each of which allows self-similar solutions: (1) the stage of initial spreading of the disc from a ring, (2) the gradual development of a quasi-stationary distribution of parameters in the disc, and (3) the late ‘spreading’ of the disc away from the center, accompanied by a decrease in the central accretion rate (see the dashed lines in Fig. 2).
@@
The disc gradually evolves into the third and final stage (the decay stage), when the details of the initial distribution are ‘forgotten’ anyway, and only some integral quantities conserved by the evolution are important in finding the self-similar solution. [^4] The nonlinear problem has the following distinctive features. Firstly, the self-similar solutions of the second kind exist only for $m\neq 0$. Secondly, self-similar solutions of the first kind in the third stage, while they exist for $m=0$, have an exponential radial profile for $r\rightarrow\infty$, characteristic for linear problems [see, for example, lyn-pri1974, and Fig. 1, the left panel]. For $m\neq 0$, the position of the outer boundary of the disc is fully determined, see the dashed profile in Fig. 2, also rafikov2016. This property is similar to the one that arises in problems of thermal conductivity, when, due to the nonlinearity, the heatwave boundary sharply separates the heated zone from the rest of the region [zeldovich-raizer1967]. The solution is sought in the form $F={h^{({n+2})/{m}}}\,{(\pm D\,t)^{-1/m}}\,y(\xi)$ where $y$ is a dimensionless function of a dimensionless argument $\xi={h}/({A\,(\pm t)^{\beta}})$. The index $\beta$ can be determined in the course of the solution for the first and second stage, and from the conservation of the angular momentum, for the decay stage (then $\beta$ can be expressed algebraically, see also pringle1974PhD). The accretion rate on the gravitating center decays as follows [see., e.g., chapter 1 in shakura_etal2018] [^5] For $a=0$ and $m=0$, Eq. (18) transforms into the lyn-pri1974 solution (10) at large $t$. :
@@
$\alpha$-disc, $\kappa=$ const, $P_{\rm tot}=P_{\rm rad}$ -2 3/2 (-8/9) 22 77
@@
$\alpha$-disc, $\kappa\propto\rho T^{-5/2}$, $P_{\rm tot}=P_{\rm gas}$ $1/2$ 11 $11/9$ $1/3$ 11 $3.319$
convective turbulence, $\kappa\propto T^{2}$, lin-bodenheimer1982 2 0 $15/14$ $2/3$ $8/3$ 4.820
molecular disc with gravitational instability, lin-pringle1987 2 $9/2$ $6/5$ $2/3$ 1/3-1/3 1.788
In the gas-pressure dominated $\alpha$-disc with the opacity according to the Kramers law and Thomson scattering, the accretion rate decreases as $\propto t^{-19/16}$ and $\propto t^{-5/4}$, respectively [pringle1974PhD, filipov1984, lyub-shak1987, cannizzo_etal1990, pringle1991]. Actually, an opacity $\kappa\propto\rho T^{-5/2}$ better describes $\alpha$-discs around stellar-mass objects for conditions near the outer radius of the hot ionized zone [tavleev+2023]. Other solutions in Table 2 are important in the context of protoplanetary discs and discs in galactic nuclei. lin-pringle1987 considered a molecular disc with a gravitational instability producing an effective viscosity. A different approach to describe analytically viscosity in self-gravitating discs, depending on their thermal structure and external irradiation, can be found in Rafikov2009, Rafikov2015. lin-bodenheimer1982 studied the evolution of a protoplanetary disc under the influence of convective turbulent viscosity.
pringle1991 considered an “external” disc with a zero central accretion rate and a finite central torque, w [a massive binary surrounded by a gas disc, see also ivanov_etal1999]. For all mass initially at the origin and being gradually expelled to infinity, the inner torque decays as
@@
The latter dependence converges to the solution for the linear equation with $a=0$, see §2.1. In rafikov2013, self-similar solutions were found with a non-zero constant mass accretion rate at the inner boundary $\dot{M}(R_{\rm in})=\chi\dot{M}_{\infty}$ and a constant external mass supply $\dot{M}_{\infty}$. Larger $\chi$ values imply less mass accumulation in the disc and a smaller inner torque. General late-time power-law asymptotics for freely expanding discs with simultaneously non-zero central accretion rate and non-zero central viscous torque at the inner boundary were presented by rafikov2016. Formulating such solutions required setting the inner coordinate of the disc to zero.
Figure 2: Distributions of viscous torque $F$ versus angular momentum $h$ at two moments of time in $\alpha$-discs, freely-expanding (dashed) and limited by the outer radius with $h_{\rm out}=1$ (solid). The accretion rate on the center decays with time.
A time-dependent solution of ogilvie1999 for accretion flows that retain the heat generated by viscous dissipation — ’advection-dominated’ flows – is another example of using similarity methods. In the case of quasi-spherical accretion, the governing equations (which are spherically averaged) are nonlinear, so that the self-similar solution of the first kind is an attractor for the initial-value problem with conserved angular momentum. The accretion rate varies $\propto t^{-4/3}$, which is consistent with the solution of the thin-disc evolution equation with z/r=z/r=const, see Table 1.
@@ 2.4 Non-linear evolution equation for $r_{\mathrm{out}}=\,$const
Figure 3: Comparison of analytic solutions (12), (16), (18), and (20) for the evolution of the accretion rate onto a 10 $M_{\odot}$ black hole with different viscosity and disc outer boundary (freely expanding or fixed). The viscosity in each solution is chosen to approximate the law in the zone of dominant gas pressure and opacity according to the Kramers formula. At late times, the time-dependencies tend to $(t/\tau_{\mathrm{pl}})^{-7/5}$, $\exp(-t/t_{\mathrm{exp}})$, $(1+t/\tilde{t}_{0})^{-5/4}$, and $(1+t/t_{0})^{-10/3}$, respectively. The input parameters are $M_{\rm disc}=6\times 10^{24}$~g, $R_{\mathrm{out}}=2\times 10^{11}$ cm and $t_{\rm vis}=60^{\mathrm{d}}$. The corresponding $e$-decay time is $t_{\mathrm{exp}}\approx 27^{\mathrm{d}}$; the bounded $\alpha$-disc’s solution is close to the linear-problem solution for about $3\,t_{\mathrm{exp}}$.
A solution to the basic equation of non-stationary accretion (3) for a disc with a constant outer radius can be found during the decay stage using the method of separation of variables: $F(h,t)=F(t)\times f(h/h_{0})$. The accretion rate changes proportionally to $F(t)$ [Ludwig+1994, lipunova_shakura2000, Ritter-King2001]:
@@
where $\dot{M}_{0}$ is the central accretion rate at $t=0$, which can be chosen as any time at the decay stage. The space part of the solution $f(h/h_{0})$ for a disc with zero (or very small) viscous stress at the inner boundary was found semi-analytically as a polynomial by lipunova_shakura2000 (see Fig. 2, the solid curves). The outer boundary condition was set to $\dot{M}(r_{\mathrm{out}})=0$, which is relevant during outbursts of X-ray novae, when the accretion rate inside the disc is much higher than the mass transfer rate from the companion star. The full solution allows determination of $t_{0}$, which is of the order of the viscous time at the disc outer radius:
@@
and the time scale on which the disc mass decays initially [Ritter-King2001]. Here $h_{\rm out}=\sqrt{GMr_{\rm out}}$, $a_{0}\equiv f^{\prime}(0)$, a=m/(1m)a=m/(1-m), and $\lambda\sim 3$ (see Table 2). The solution for Kramers opacity was used to model the optical and X-ray lightcurves of the X-ray novae A 0620-00 and GU Mus 1124-68 during the decline after the peak of their outbursts and to obtain constraints on the turbulence parameter $\alpha$ [lip-sha2002, suleimanov_etal2008].